Two horizontal parallel plates are conducting in a vacuum.

A small particle of mass 4 × 10-12kg, carrying a positive charge of 3.0 × 10-14C is released at A close to the upper plate. What total force acts on this particle?


1
Expert's answer
2021-01-10T18:31:15-0500

Explanation & Calculations


  • Since the plates are conducting, they generate a magnetic field around them.
  • Plates should be conducting in opposite directions in order to maintain an effective magnetic field between the plates otherwise they cancel out.
  • Consider the figure attached



  • "\small B = \frac{1}{2}\mu_0J" : "\small j=" current per unit length of the plate
  • This field generates a force (F) on the charge as shown (according to the left-hand rule)

"\qquad\qquad \small F=2Bqv"

  • As the particle is released from rest, F=0 at the beginning & increases as it falls down under gravity.
  • Therefore, the net force which continues to act is

"\qquad\qquad
\begin{aligned}
\small F_{net}&= \small \sqrt{mg^2+F^2}\\
&= \small \sqrt{(mg)^2+(2Bqv)^2}
\end{aligned}"

  • This moves the particle in a trajectory.